OpenStudy (anonymous):

Solve on the interval [0,2π): 1+sinθ=1/2 A. 7π/6, 11π/6 B. 2π/3, 4π/3 C. π/3, 5π/3 D. π/6, 5π/6

OpenStudy (anonymous):

sina = -1/2 option A

OpenStudy (anonymous):

sin theta =-1/2 theta=npi+- pi/6 theta=7pi/6 and 11pi/6